# Rigidity for Patterson–Sullivan systems with applications to random walks and entropy rigidity

### v2

@@ Abstract.
In this paper we introduce Patterson–Sullivan systems, which consist of a group action on a compact metrizable space and a quasi-invariant measure which behaves like a classical Patterson–Sullivan measure. For such systems we prove a generalization of Tukia’s measurable boundary rigidity theorem. We then apply this generalization to (1) study the singularity conjecture for Patterson–Sullivan measures (or, conformal densities) and stationary measures of random walks on isometry groups of Gromov hyperbolic spaces, mapping class groups, and discrete subgroups of semisimple Lie groups; (2) prove versions of Tukia’s theorem for word hyperbolic groups, Teichmüller spaces, and higher rank symmetric spaces; and (3) prove an entropy rigidity result for pseudo-Riemannian hyperbolic spaces.
@@ 1.1.1. Random walks on Gromov hyperbolic spaces
[…]
@@ 1.1.2. Random walks on mapping class groups and Teichmüller spaces
We also note that Athreya–Bufetov–Eskin–Mirzakhani [ABEM12] constructed a Patterson–Sullivan measure for $\operatorname{Mod}(\Sigma)$ on $\mathcal{PMF}$ using Thurston measure, and Gekhtman [Gek12] constructed Patterson–Sullivan measures for convex cocompact subgroups of $\operatorname{Mod}(\Sigma)$ on $\mathcal{UE}$. Since the identity map $\operatorname{\mathcal{T}}\to\operatorname{\mathcal{T}}$ continuously extends to a topological embedding $\mathcal{UE}\hookrightarrow\partial_{GM}\operatorname{\mathcal{T}}$ [Miy13], the Patterson–Sullivan measures constructed in [Gek12] are Patterson–Sullivan measures on $\partial_{GM}\operatorname{\mathcal{T}}$. Further, by works of Masur [Mas82] and Veech [Vee82], the Patterson–Sullivan measure constructed in [ABEM12] gives a full measure on $\mathcal{UE}$, and therefore can be identified with a Busemann Patterson–Sullivan measure on $\partial_{GM}\operatorname{\mathcal{T}}$.
@@
There are many examples of subgroups of $\operatorname{Mod}(\Sigma)$ which are relatively hyperbolic and containing multitwists, so Theorem 1.11 applies to. For instance, the combination theorem for Veech subgroups by Leininger–Reid [LR06] produces closed surface subgroups in $\operatorname{Mod}(\Sigma)$ with multitwists, and so-called parabolically geometrically finite subgroups introduced by Dowdall–Durham–Leininger–Sisto [DDLS24] are relatively hyperbolic and contain multitwists in their peripheral subgroups. Many examples of parabolically geometrically finite subgroups were also constructed by Udall [Uda25] and Loa [Loa21]. Finally, in their proof of the purely pseudo-Anosov surface subgroup conjecture, Kent–Leininger [KL24] constructed a type-preserving homomorphism from a finite index subgroup of the fundamental group of the figure-8 knot complement into $\operatorname{Mod}(\Sigma)$ when $\Sigma$ has genus at least 4. The image of such a homomorphism is relatively hyperbolic and contains a multitwist.
@@ 1.1.3. Random walks on discrete subgroups of Lie groups
Let $\operatorname{\mathsf{G}}$ be a connected semisimple Lie group without compact factors and with finite center. Suppose $\Gamma<\operatorname{\mathsf{G}}$ is a Zariski dense discrete subgroup, and $\mathsf{m}$ is a probability measure on $\Gamma$ whose support generates $\Gamma$ as a semigroup. Fix a minimal parabolic subgroup $\operatorname{\mathsf{P}}$ and let $\operatorname{\mathcal{F}}:=\operatorname{\mathsf{G}}/\operatorname{\mathsf{P}}$ denote the Furstenberg boundary. Then there is a unique $\mathsf{m}$-stationary measure $\nu$ on $\operatorname{\mathcal{F}}$ [Fur73, GR85]. The measure $\nu$ is also referred to as the Furstenberg measure.
In this section we consider the following well-known conjecture (cf. Kaimanovich–Le Prince [KLP11]).
@@
If $\mathsf{m}$ has finite support, then the $\mathsf{m}$-stationary measure $\nu$ is singular to the Lebesgue measure class on $\operatorname{\mathcal{F}}$.
@@
If $\mathsf{m}$ has finite support, then the $\mathsf{m}$-stationary measure $\nu$ is singular to every coarse Iwasawa Patterson–Sullivan measure on $\operatorname{\mathcal{F}}$.
@@
Suppose $\Gamma$ is relatively hyperbolic (as an abstract group), $\mathsf{m}$ has finite superexponential moment, and $\mu$ is a coarse $\phi$-Patterson–Sullivan measure on $\operatorname{\mathcal{F}}$ of dimension $\delta$. If the measures $\nu$ and $\mu$ are not singular, then:
@@
    $\sup_{\gamma\in\Gamma}\left|\operatorname{d}_{G}(\operatorname{id},\gamma)-\delta\phi(\kappa(\gamma))\right|<+\infty$. In particular, $\sum_{\gamma\in\Gamma}e^{-\delta\phi(\kappa(\gamma))}=~{}+\infty$ and $\delta\phi\in\mathfrak{a}^{*}$ is tangent to the growth indicator of $\Gamma$.
@@
Suppose $\Gamma$ is word hyperbolic (as an abstract group) and $\mathsf{m}$ has finite superexponential moment. If $\Gamma$ contains a unipotent element of $\operatorname{\mathsf{G}}$, then the $\mathsf{m}$-stationary measure $\nu$ is singular to every coarse Iwasawa Patterson–Sullivan measure on $\operatorname{\mathcal{F}}$.
@@ 1.2.2. Tukia’s theorem for Teichmüller spaces
  • There exists an onto homomorphism $\rho:\Gamma_{1}\to\Gamma_{2}$ and a $\mu_{1}$-almost everywhere defined measurable $\rho$-equivariant injective map $f:\partial_{GM}\operatorname{\mathcal{T}}_{1}\to~{}\partial_{GM}\operatorname{\mathcal{T}}_{2}$.
@@ 1.2.3. Tukia’s theorem in higher rank
1.3. Entropy rigidity in pseudo-Riemannian hyperbolic geometry
Delaying more definitions until Section 13, let $\operatorname{\mathbb{H}}^{p,q}$ be pseudo-Riemannian hyperbolic space of signature $(p,q)$. The group $\mathsf{PO}(p,q+1)$ acts by isometries on this pseudo-metric space and using this action Danciger–Guéritaud–Kassel [DGK18] introduced $\operatorname{\mathbb{H}}^{p,q}$-convex cocompact subgroups of $\mathsf{PO}(p,q+1)$. Glorieux–Monclair [GM21] introduced a critical exponent $\delta_{\operatorname{\mathbb{H}}^{p,q}}(\Gamma)$ for a convex cocompact subgroup $\Gamma<\mathsf{PO}(p,q)$ and proved that
$\delta_{\operatorname{\mathbb{H}}^{p,q}}(\Gamma)\leq p-1.$
The critical exponent $\delta_{\operatorname{\mathbb{H}}^{p,q}}(\Gamma)$ is also referred to as entropy of $\Gamma$.
Using our version of Tukia’s theorem for higher rank Lie groups (Theorem 1.21), we characterize the equality case.
Theorem 1.24 (see Theorem 13.2 below).
If $\Gamma<\mathsf{PO}(p,q+1)$ is $\operatorname{\mathbb{H}}^{p,q}$-convex cocompact and $\delta_{\operatorname{\mathbb{H}}^{p,q}}(\Gamma)=p-1$, then $\Gamma$ preserves and acts cocompactly on a totally geodesic copy of $\operatorname{\mathbb{H}}^{p}$ in $\operatorname{\mathbb{H}}^{p,q}$.
Remark 1.25.
A totally geodesic copy of $\operatorname{\mathbb{H}}^{k}$ in $\operatorname{\mathbb{H}}^{p,q}$ is a subset of the form $\operatorname{\mathbb{P}}(V)\cap\operatorname{\mathbb{H}}^{p,q}$ where $V\subset\operatorname{\mathbb{R}}^{p+q+1}$ is a $(k+1)$-dimensional linear subspace and the associated bilinear form $[\cdot,\cdot]_{p,q+1}$ restricted to $V$ has signature $(k,1)$.
A number of special cases of Theorem 1.24 have been previously established:
  1. (1)
    $\operatorname{\mathbb{H}}^{p,0}$ is real hyperbolic $p$-space and $\operatorname{\mathbb{H}}^{p,0}$-convex cocompact coincides with the usual definition in real hyperbolic geometry. In this case, the above theorem follows from a result of Tukia [Tuk84], which also shows that a non-lattice geometrically finite group has critical exponent strictly less than $p-1$.
  2. (2)
    Collier–Tholozan–Toulisse [CTT19] proved the above theorem when $p=2$ and $\Gamma$ is the fundamental group of a closed surface.
  3. (3)
    Mazzoli–Viaggi [MV23] proved the above theorem when $\Gamma$ is the fundamental group of a closed $p$-manifold.
The techniques used in [CTT19, MV23] strongly use the fact that $\Gamma$ is the fundamental group of a closed manifold and are very different than the approach taken here. In the proof of Theorem 1.24 we construct coarse Iwasawa Patterson–Sullivan measures on two different flag manifolds and show that there is a measurable map so that the push-forward of one of the measures is non-singular to the other. Then we use Theorem 1.21 to constrain the eigenvalues of elements in the group, which in turn constrains the Zariski closure of the group.
@@ A Patterson–Sullivan-system (PS-system) of dimension $\delta$ consists of
    If $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\rightarrow+\infty$, $Z\subset M$ is compact, and $[M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance, then for any $x\in Z$, there exists $g\in\Gamma$ such that
@@
:=$\mathscr{H}:=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$
@@
    If $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\rightarrow+\infty$, $Z\subset M$ is compact, and $[M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance, then for any $h_{1},\ldots,h_{m}\in\Gamma$ and $x\in Z$, there exists $g\in\Gamma$ such that
@@
For a well-behaved Patterson–Sullivan system $(M,\Gamma,\sigma,\mu)$ with respect to a hierarchy =$\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$, we consider the following analogue of the conical limit set:
@@
    Coarse Iwasawa PS-measures on a partial flag manifold associated to Zariski dense subgroups (more generally “$\operatorname{\mathsf{P}}_{\theta}$-irreducible” subgroups) are always contained in PS-systems. When the subgroup is transverse and the measure is supported on the limit set, they are contained in well-behaved PS-systems (see Section 9; see also Theorem 9.12 for general Zariski dense discrete subgroups).
@@ Acknowledgements
We would like to thank Dick Canary, Sebastian Hurtado-Salazar, Yair Minsky, and Hee Oh for valuable conversation. Kim expresses his special gratitude to his Ph.D. advisor Hee Oh for her encouragement and guidance.
@@ 2.1. Possibly ambiguous notation/terminology
    A sequence $\{y_{n}\}$ in a countable set $Y$ is escaping if it eventually leaves every finite set, i.e. if $F\subset Y$ is finite, then #$\#\{n:y_{n}\in F\}$ is finite.
@@ 3. Basic properties of PS-systems
[…]
@@
Let $(M,\Gamma,\sigma,\mu)$ be a well-behaved PS-system with respect to a hierarchy =$\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$. Let $R>0$ and let $R^{\prime}>0$ be the constant satisfying Property (PS7) for $R$. Then for any $I\subset\mathscr{H}(R)$, there exists $J\subset I$ such that
@@ 4. An analogue of the conical limit set
Using this notation, the conical limit set of a hierarchy =$\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$ defined in Equation (9) can be rewritten as
@@ 5. An analogue of the Lebesgue differentiation theorem
Let $(M,\Gamma,\sigma,\mu)$ be a well-behaved PS-system of dimension $\delta\geq 0$ with respect to a hierarchy =$\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$. Fix $R_{0}>0$ such that any $R\geq R_{0}$ satisfies the Shadow Lemma (Proposition 3.1).
@@
[…]
@@
[…]
@@
$N_{1}$ :=$:=\{x:\mathcal{B}_{\alpha}(h-g)(x)>c/3\};$
$N_{2}$ :=$:=\{x:\left|h(x)-g(x)\right|>c/3\};$
$N_{3}$ :=$:=\{x:\mathcal{A}_{\alpha}g(x)>c/3\}.$
@@
[…]
@@
[…]
@@
[…]
@@
00 0 $\leq\lim_{n\rightarrow\infty}\frac{1}{\mu(\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}))}\mu\left(\left\{y\in\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}):\operatorname{d}_{Y}(F(x),F(y))>\epsilon\right\}\right)$
[…]

### v3

@@ Abstract.
In this paper we introduce Patterson–Sullivan systems, which consist of a group action on a compact metrizable space and a quasi-invariant measure which behaves like a classical Patterson–Sullivan measure. For such systems we prove a generalization of Tukia’s measurable boundary rigidity theorem. We then apply this generalization to (1) study the singularity conjecture for Patterson–Sullivan measures (or, conformal densities) and stationary measures of random walks on isometry groups of Gromov hyperbolic spaces, mapping class groups, and discrete subgroups of semisimple Lie groups; (2) prove versions of Tukia’s theorem for word hyperbolic groups, Teichmüller spaces, and higher rank symmetric spaces; and (3) in a companion paper prove an entropy rigidity result for Anosov groups with Lipschitz limit sets.
@@
Patterson–Sullivan measures, random walks, Gromov hyperbolic metric spaces, mapping class groups, semisimple Lie groups, entropy rigidity
2020 Mathematics Subject Classification:
20H10, 37F32, 37H12, 22E40, 37A99, 57K20, 30F60
@@ 1.1.1. Random walks on Gromov hyperbolic spaces
for any $c>1$, where $\left|\cdot\right|$ is the distance from the identity with respect to a word metric on $\Gamma$.
@@ 1.1.2. Random walks on mapping class groups and Teichmüller spaces
We also note that Athreya–Bufetov–Eskin–Mirzakhani [ABEM12] constructed a Patterson–Sullivan measure for $\operatorname{Mod}(\Sigma)$ on $\mathcal{PMF}$ using Thurston measure, and Gekhtman [Gek12] constructed Patterson–Sullivan measures for convex cocompact subgroups of $\operatorname{Mod}(\Sigma)$ on $\mathcal{UE}$. Since the identity map $\operatorname{\mathcal{T}}\to\operatorname{\mathcal{T}}$ continuously extends to a topological embedding $\mathcal{UE}\hookrightarrow\partial_{GM}\operatorname{\mathcal{T}}$ [Miy13], the Patterson–Sullivan measures in [Gek12] are Patterson–Sullivan measures on $\partial_{GM}\operatorname{\mathcal{T}}$. Further, by works of Masur [Mas82] and Veech [Vee82], the Patterson–Sullivan measure constructed in [ABEM12] gives a full measure on $\mathcal{UE}$, and therefore can be identified with a Busemann Patterson–Sullivan measure on $\partial_{GM}\operatorname{\mathcal{T}}$.
@@
There are many examples of subgroups of $\operatorname{Mod}(\Sigma)$ which are relatively hyperbolic and containing multitwists, so Theorem 1.11 applies to. For instance, the combination theorem for Veech subgroups by Leininger–Reid [LR06] produces closed surface subgroups in $\operatorname{Mod}(\Sigma)$ with multitwists, and so-called parabolically geometrically finite subgroups introduced by Dowdall–Durham–Leininger–Sisto [DDLS24] are relatively hyperbolic and contain multitwists in their peripheral subgroups. Many examples of parabolically geometrically finite subgroups were also constructed by Udall [Uda25], Aougab et al. [ABD+25], and Loa [Loa21]. Finally, in their proof of the purely pseudo-Anosov surface subgroup conjecture, Kent–Leininger [KL24] constructed a type-preserving homomorphism from a finite index subgroup of the fundamental group of the figure-8 knot complement into $\operatorname{Mod}(\Sigma)$ when $\Sigma$ has genus at least 4. The image of such a homomorphism is relatively hyperbolic and contains a multitwist.
@@ 1.1.3. Random walks on discrete subgroups of Lie groups
[…]
In this section we consider the following well-known conjecture, cf. [KLP11].
@@
[…]
@@
[…]
@@
[…]
@@
    $\sup_{\gamma\in\Gamma}\left|\operatorname{d}_{G}(\operatorname{id},\gamma)-\delta\phi(\kappa(\gamma))\right|<+\infty$. In particular,
    $\sum_{\gamma\in\Gamma}e^{-\delta\phi(\kappa(\gamma))}=~+\infty$
    and $\delta\phi\in\mathfrak{a}^{*}$ is tangent to the growth indicator of $\Gamma$.
@@
[…]
@@ 1.2.2. Tukia’s theorem for Teichmüller spaces
  • There exists an onto homomorphism $\rho:\Gamma_{1}\to\Gamma_{2}$ and a $\mu_{1}$-almost everywhere defined measurable $\rho$-equivariant injective map $f:\partial_{GM}\operatorname{\mathcal{T}}_{1}\to\partial_{GM}\operatorname{\mathcal{T}}_{2}$.
@@ 1.2.3. Tukia’s theorem in higher rank
1.3. Entropy rigidity for Anosov groups with Lipschitz limit sets
The Hilbert critical exponent of a discrete subgroup $\Gamma<\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is
$\delta_{\rm Hil}(\Gamma):=\limsup_{R\rightarrow\infty}\frac{1}{R}\log\#\left\{\gamma\in\Gamma:\frac{1}{2}\log\left\|\gamma\right\|\left\|\gamma^{-1}\right\|\leq R\right\}.$
Notice that if $\Gamma$ preserves a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d})$, then $\delta_{\rm Hil}(\Gamma)$ coincides with the critical exponent of $\Gamma$ with respect to the Hilbert metric on $\Omega$ (this follows from [DGK24, Proposition 10.1]).
In a companion paper, we use Corollary 9.14 (the general version of Theorem 1.21) to prove a rigidity result for this quantity for Anosov groups with Lipschitz limit sets. We briefly and informally describe this result here, for more details see  [KZ26]. Pozzetti–Sambarino–Wienhard [PSW23] proved that
$\delta_{\rm Hil}(\Gamma)\leq p-1$
when $\Gamma<\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is projective Anosov whose projective limit set in $\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d})$ is a $p$-dimensional Lipschitz manifold and $\Gamma$ acts irreducibly on $\operatorname{\mathbb{R}}^{d}$ and $\wedge^{p+1}\operatorname{\mathbb{R}}^{d}$, for some $p\leq d-2$. Under this hypothesis, in [KZ26, Theorem 1.13] we prove that
$\delta_{\rm Hil}(\Gamma)=p-1$
if and only if p=d2p=d-2 and up to conjugation $\Gamma$ is a uniform lattice in $\mathsf{SO}(d-1,1)$.
This generalizes a result of Crampon [Cra09] who proved that if $\Gamma<\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ is a discrete group which acts cocompactly on a strictly convex domain $\Omega\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d})$, then
$\delta_{\rm Hil}(\Gamma)\leq d-2$
with equality if and only if $\Omega$ is an ellipsoid (and hence up to conjugation $\Gamma$ is a uniform lattice in $\mathsf{SO}(d-1,1)$). Under Crampon’s hypothesis, $\Gamma$ is $\operatorname{\mathsf{P}}_{1}$-Anosov and acts irreducibly on $\operatorname{\mathbb{R}}^{d}$ and hence also $\wedge^{d-1}\operatorname{\mathbb{R}}^{d}$, see [GW12, Section 6.2]. Further, $\Lambda_{1}(\Gamma)$ coincides with $\partial\Omega$ and hence is a Lipschitz (d2)(d-2)-manifold.
In the proof of [KZ26, Theorem 1.13] we construct a measurable boundary map from the projective limit set into the partial flag manifold $\operatorname{\mathcal{F}}_{1,p+1,d-p-1,d-1}(\operatorname{\mathbb{R}}^{d})$ and then show that two Patterson–Sullivan measures are non-singular. Then Corollary 9.14 constrains the Jordan projection and we use a result of Benoist [Ben97] to determine the Zariski closure.
@@ A Patterson–Sullivan-system (PS-system) of dimension $\delta$ consists of
    If $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\rightarrow+\infty$, $Z\subset M$ is compact, and $M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance, then for any $x\in Z$, there exists $g\in\Gamma$ such that
@@
$\mathscr{H}:=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$
@@
    If $\{\gamma_{n}\}\subset\Gamma$, $R_{n}\rightarrow+\infty$, $Z\subset M$ is compact, and $M\smallsetminus\gamma_{n}^{-1}\operatorname{\mathcal{O}}_{R_{n}}(\gamma_{n})]\rightarrow Z$ with respect to the Hausdorff distance, then for any $h_{1},\ldots,h_{m}\in\Gamma$ and $x\in Z$, there exists $g\in\Gamma$ such that
@@
For a well-behaved Patterson–Sullivan system $(M,\Gamma,\sigma,\mu)$ with respect to a hierarchy $\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$, we define the $\mathscr{H}$-conical limit set:
@@
    Coarse Iwasawa PS-measures on a partial flag manifold associated to Zariski dense subgroups (more generally $\operatorname{\mathsf{P}}_{\theta}$-irreducible” subgroups) are always contained in PS-systems. When the subgroup is transverse and the measure is supported on the limit set, they are contained in well-behaved PS-systems (see Section 9; see also Theorem 9.12 for general Zariski dense discrete subgroups).
@@ Acknowledgements
We would like to thank Harry Bray, Dick Canary, Sebastian Hurtado-Salazar, Yair Minsky, and Hee Oh for valuable conversation. Kim expresses his special gratitude to his Ph.D. advisor Hee Oh for her encouragement and guidance.
@@ 2.1. Possibly ambiguous notation/terminology
    A sequence $\{y_{n}\}$ in a countable set $Y$ is escaping if it eventually leaves every finite set, i.e. if $F\subset Y$ is finite, then $\{n:y_{n}\in F\}$ is finite.
@@ 3. Basic properties of PS-systems
$\mu(\operatorname{\mathcal{O}}_{R}(\gamma))=\int_{\gamma^{-1}\operatorname{\mathcal{O}}_{R}(\gamma)}\frac{d\gamma^{-1}_{*}\mu}{d\mu}d\mu,$
@@
Let $(M,\Gamma,\sigma,\mu)$ be a well-behaved PS-system with respect to a hierarchy $\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$. Let $R>0$ and let $R^{\prime}>0$ be the constant satisfying Property (PS7) for $R$. Then for any $I\subset\mathscr{H}(R)$, there exists $J\subset I$ such that
@@ 4. An analogue of the conical limit set
Using this notation, the $\mathscr{H}$-conical limit set defined in Equation (9) can be rewritten as
@@ 5. An analogue of the Lebesgue differentiation theorem
Let $(M,\Gamma,\sigma,\mu)$ be a well-behaved PS-system of dimension $\delta\geq 0$ with respect to a hierarchy $\mathscr{H}=\{\mathscr{H}(R)\subset\Gamma:R\geq 0\}$. Fix $R_{0}>0$ such that any $R\geq R_{0}$ satisfies the Shadow Lemma (Proposition 3.1).
@@
$0=\lim_{n\rightarrow\infty}\frac{1}{\mu(\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}))}\int_{\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n})}|h(y)-h(x)|d\mu(y),$
@@
$h(x)=\lim_{n\rightarrow\infty}\frac{1}{\mu(\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}))}\int_{\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n})}h(y)d\mu(y),$
@@
$N_{1}$ $:=\{x:\mathcal{B}_{\alpha}(h-g)(x)>c/3\};$
$N_{2}$ $:=\{x:\left|h(x)-g(x)\right|>c/3\};$
$N_{3}$ $:=\{x:\mathcal{A}_{\alpha}g(x)>c/3\}.$
@@
$\int_{\alpha\operatorname{\mathcal{O}}_{R}(\gamma_{x})}\left|h-g\right|d\mu>\frac{c}{4}\mu(\alpha\operatorname{\mathcal{O}}_{R}(\gamma_{x})).$
@@
$\limsup_{n\to\infty}\frac{1}{\mu(\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}))}\int_{\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n})}\left|h(y)-h(x)\right|d\mu(y)\leq\mathcal{A}_{\gamma}h(x)=0,$
@@
$Ce^{-\delta\left\|\gamma_{n}\right\|}\leq\frac{d{\gamma_{n}^{-1}}_{*}\mu}{d\mu}\leq Ce^{-\delta\left\|\gamma_{n}\right\|}\quad\mu\text{-a.e.}$
@@
$0$ $\leq\lim_{n\rightarrow\infty}\frac{1}{\mu(\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}))}\mu\left(\left\{y\in\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}):\operatorname{d}_{Y}(F(x),F(y))>\epsilon\right\}\right)$
$\leq\lim_{n\rightarrow\infty}\frac{1}{\mu(\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n}))}\int_{\gamma\operatorname{\mathcal{O}}_{R}(\gamma_{n})}\left|h_{k,K}(x)-h_{k,K}(y)\right|d\mu(y)=0.$